# EZIGen: Enhancing zero-shot personalized image generation with precise subject encoding and decoupled guidance

### v3

@@ 4.5 Comparing feature extractor with other choices
  ID Preser. Prompt Consis. Overall Qaulity
21.7% 20.6% 20.3%
36.5% 45.2% 37.5%
41.8% 34.1% 42.2%
Table 3: Left: Quantitative evaluation following [40]. Right: Human preferences for each metric and the overall generation quality. EZIGen shows very competitive results in domain-specific generation tasks WITHOUT using domain-specific datasets or large-scale pre-training, it also achieves a perfect balance between ID preservation and prompt consistency. These results further highlight our design, reaching high in both aspects without sacrifices.
@@ 4.6 Zero-shot abilities for domain-specific task
To demonstrate the versatilities of our model, we further evaluate the effectiveness of our method for a domain-specific generation task, namely personalized human image generation. Compared to generating common objects, producing accurate human faces is extremely challenging as human brains possess strong overfittings to human facial landmarks, making even the slightest distortion prominent. We report the performance following FastComposer’s evaluation protocols and report the results in Tab. 3 and Fig. 17 in the appendix. Unlike FastComposer[40], which relies on domain-specific datasets, or Subject Diffusion [28], which requires large-scale pretraining, our model achieves high-quality results using only a normal-scale open-domain dataset. As shown in Tab. 3, our approach delivers very competitive results. Regarding subject ID preservation, we match Subject Diffusion and outperform FastComposer. While our prompt consistency lags behind FastComposer and DreamBooth, it still exceeds Subject Diffusion. These findings further prove the capabilities of EZIGen in capturing fine-grained subject details regardless of categories.
@@ 5 Conclusion
Appendix A Dataset construction
@@
YoutubeVIS Dataset. The YoutubeVIS dataset contains videos with annotated instances of objects over time. To create training pairs, we extract images of the same subject from different frames of a video, following the methodology proposed in [7]. This process captures the variations in appearance, pose, and position of the same subject across different frames, providing valuable temporal data that helps the model learn consistent subject identification even in dynamic scenes.
  Method CLIP-T DINO # Sub TS DS ZS
Textual Inversion 0.261 0.561 3-6 N/A N/A
DreamBooth 0.306 0.672 4-6 N/A N/A
Elite 0.296 0.647 1 multi 0.125M
BLIP-Diffusion 0.298 0.589 1 multi $\sim$2M
IP-Adapter 0.274 0.608 1 single 10M
SSR-Encoder 0.308 0.612 1 single 10M
BootPIG 0.311 0.674 4-6 single 0.2M
Subject Diffusion 0.293 0.711 1 single $\sim$76M
Ours 0.316 0.718 1 single 0.2M
Table 5: More quantitative comparisons on DreamBench. For each abbreviation, “#Sub” stands for the number of subject images required, and “TS” indicates if the methods are trained end-to-end or require multiple stages of pertaining. “DS” represents the dataset size required for achieving the reported scores, and “ZS” shows if the method is capable of zero-shot inference. Our method achieves the most satisfying results with the least system complexity compared to previous methods.
@@
Figure 16: Integration with stronger base image models, namely Flux and Stable Diffusion 3.
Appendix B Autostop machanism
@@
Appendix C Justify evaluation metric: CLIP-I v.s. DINO
@@
Appendix D User study details
Except for quantitative comparisons, we further evaluate our methods based on human preferences. These results are listed in Tab. 1, Tab. 2 and Tab. 3. Specifically, for each evaluation metric, human evaluators select their preferred images the best match indicated evaluation metric from batches of shuffled images generated by each competing method. The preferences are then aggregated across batches to calculate the percentages reported in each table.
Appendix E Integrating with existing T2I generators
E.1 Training free integration
Despite the fact that the method is developed based on Stable Diffusion 2.1 base, it could be naturally integrated with existing state-of-the-art base models. For instance, in Fig. 16, we regard our EZIGen model as a subject-driven image editor and replace the specific concepts within the original image generated by advanced image models such as Stable Diffusion 3 and Flux models.
E.2 Training on advanced base models
We also empirically presume that our concept and designs are suitable for tuning all diffusion- or flow-matching-based models as they all share a similar add-noise/denoising process as Stable Diffusion 2.1, therefore, we believe that with proper datasets and base-model-related structure modifications, the trained models are expected to obtain similar or even better results compared to the current design.
Appendix F Limitations and future work
EZIGen may sometimes fail to generate images containing subjects with drastic pose variations, for example, given an image of a dog facing toward the camera, it would be rather challenging to generate an image with the dog facing forward. Moreover, our model inherits the weakness from Stable Diffusion 2.1 that the model cannot output accurate detailed text characters. Future possibilities for a more advanced model include migrating the subject encoding technique to more advanced image/video base models.
Figure 18: More visualization results for personalized image generation.

### v4

@@ 4.5 Comparing feature extractor with other choices
Ours+SDXL 0.601 0.246
  ID Prompt Overall
Preser. Consis. Quality
0.13 0.20 0.07
0.27 0.32 0.28
0.21 0.35 0.35
0.39 0.13 0.30
Table 3: Left: Quantitative evaluation following [45]. Right: Human preferences Since most of the comparable methods are based on relatively older models, we choose to perform human evaluation on our SD2.1-base implementation. For each metric and the overall quality. EZIGen shows very competitive results in domain-specific generation tasks WITHOUT using domain-specific datasets or large-scale pre-training, it also achieves a perfect balance between ID preservation and prompt consistency. These results further highlight our design, reaching high in both aspects without sacrifices.
@@ 4.6 Zero-shot abilities for domain-specific task
To demonstrate the versatility of our model, we further evaluate the effectiveness of our method for a domain-specific generation task, namely personalized human image generation. Compared to generating common objects, producing accurate human faces is extremely challenging as human brains possess strong overfitting to human facial landmarks, making even the slightest distortion prominent. We report the performance following FastComposer’s evaluation protocols and report the results in Tab. 3 and Fig. 16 in the appendix. Unlike FastComposer[45], which relies on domain-specific datasets, or Subject Diffusion [31], which requires large-scale pretraining, our model achieves high-quality results using only a normal-scale open-domain dataset. As shown in Tab. 3, our approach delivers very competitive results. Regarding subject ID preservation, we match Subject Diffusion and outperform FastComposer. While our prompt consistency lags behind FastComposer and DreamBooth, it still exceeds Subject Diffusion. These findings further prove the capabilities of EZIGen in capturing fine-grained subject details regardless of categories.
@@ 5 Conclusion
6 User study details
Except for quantitative comparisons, we further evaluate our methods based on human preferences. These results are listed in Tab. 1, Tab. 2 and Tab. 3. Specifically, for each evaluation metric, human evaluators select their preferred images that best match the indicated evaluation metric from the batches of shuffled images generated by each competing method. The preferences are then aggregated across batches to calculate the percentages (range from 0-1) reported in each table. For all three benchmarks, we generate 100 images for each method using random subjects and text prompts from each dataset. We then present the evaluator batch by batch, where each batch contains a subject, a text prompt, and the shuffled generated images from all methods to be evaluated. We then ask the evaluator to decide which one matches the criteria, the most (e.g., Which one best preserves the subject ID, or, personally which one do you think owns the highest overall quality), then we could easily calculate the numbers each methods win and obtain the human evaluation results. In total, 10 people from other majors, research fields participated in the experiments.
7 Dataset construction
@@
YoutubeVIS Dataset. The YoutubeVIS dataset contains videos with annotated instances of objects over time. To create training pairs, we extract images of the same subject from different frames of a video, following the methodology proposed in [8]. This process captures the variations in appearance, pose, and position of the same subject across different frames, providing valuable temporal data that helps the model learn consistent subject identification even in dynamic scenes. Subject200k Dataset. We directly adopt the images from Subject200k dataset [40] to train our SDXL model as a subset. It provides abundant high-res data to preserve the high-quality generation abilities of SDXL.
@@
Figure 14: Integration with stronger base image models, namely Flux and Stable Diffusion 3, using EZIGen as an external editor.
Figure 15: Extra editing results for “A man riding a horse”.
8 Autostop machanism
@@
9 Justify evaluation metric: CLIP-I v.s. DINO
@@
10 Integrating with any existing T2I generators
10.1 Training based integration
As already introduced in the main passage, we apply EZIGen on two base models, namely Stable Diffusion2.1 and SDXL, this demonstrates the strong applicabilities of EZIGen for generalizing to other structures.
10.2 Training free integration
In the training free setting, our model can be regarded as an off-the-shelf editor that could be naturally integrated with existing state-of-the-art base models. For instance, in Fig. 15, we regard our EZIGen model as a subject-driven image editor and replace the specific concepts within the original image generated by advanced image models such as Stable Diffusion 3 and Flux models.
11 Limitations and future work
EZIGen may sometimes fail to generate images containing subjects with drastic pose variations, for example, given an image of a dog facing toward the camera, it would be rather challenging to generate an image with the dog facing forward. Moreover, our model inherits the weakness from Stable Diffusion2.1/XL that the model cannot output accurate detailed text characters. Future possibilities for a more advanced model include migrating the subject encoding technique to more advanced image/video-based models.
Figure 17: More visualization results for personalized image generation using Stable Diffusion XL.
Figure 18: More visualization results for personalized image generation using Stable Diffusion 2.1-base.